Synovec Company has a debt–equity ratio of .85. Return on assets is 9.6 percent, and total equity is $805,000.

What is the company's equity multiplier? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

  Equity multiplier   

What is the company's return on equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  Return on equity %
What is the company's net income? (Do not round intermediate calculations.)
  Net income $   
Answer #1

debt–equity ratio =debt/equity

Hence debt=0.85*805,000=$684250

Total assets=debt+equity

=684250+805,000=$1489250

Equity multiplier=Total assets/equity

=1489250/805,000

=1.85

ROA=net income/Total assets

net income=1489250*9.6%=$142968

ROE=net income/equity

=142968/805,000=17.76%